### v3

@@ Abstract
[Abstract]Financial reinforcement learning (FinRL) has emerged as a promising paradigm for sequential decision-making in financial engineering. However, applying RL in real-world trading tasks remains challenging due to the non-stationarity of financial data, low signal-to-noise ratios, and various market frictions. Although numerous FinRL methods have been developed for tasks such as trading and portfolio management, the lack of standardized task definitions, datasets, environments, and baselines has hindered consistent evaluation and reproducibility. To bridge this gap, we organized three FinRL Contests from 2023 to 2025, covering a diverse range of financial tasks such as stock trading, order execution, cryptocurrency trading, and the use of large language model (LLM)-generated signals. These contests attracted 200 participants from over 100 institutions across 22 countries. To promote reproduction, we provided open-source starter kits featuring GPU-optimized parallel market environments and comprehensive documentation. In this paper, we summarize these benchmarking efforts, detailing task formulations, data curation pipelines, environment implementations, evaluation protocols, participant performance, and key organizational insights.
@@
Financial reinforcement learning, FinRL, LLM, benchmark, financial engineering
articletype: Article Typejournal: Journalvolume: $0$
@@ 1 Introduction
Financial reinforcement learning (FinRL) 1, 2, 3, 4, 5, 6, 7 is an interdisciplinary field that applies reinforcement learning algorithms to financial tasks, such as algorithmic trading 8, portfolio management gu2023portfolio, option pricing vittori2021option, hedging, and market making. FinRL aims to train trading agents to interact with dynamic financial environments and to optimize their financial decisions autonomously. As FinRL methods continue to advance, two practical challenges have become increasingly apparent. One is policy instability, where small changes in training settings or environments can lead to significant variation in performance. The other is the sampling bottleneck, as collecting high-quality financial trajectories is often expensive and limited by data access. Additionally, it remains difficult to compare the performance of these RL strategies due to inconsistencies in task definitions, dataset choices, environment implementations, and baselines. These challenges make it difficult to obtain reliable and consistent results, reinforcing the need for reproducible benchmarks that enable fair and systematic evaluation across tasks.
@@
FinRL-Meta 9, 10 partially addressed this issue by transforming financial data into gym-style market environments, which built an automated data curation pipeline to deal with dynamic market data. This framework enables different FinRL algorithms to be implemented and evaluated under standardized conditions across various tasks. Still, there is no widely accepted benchmark for evaluating RL agents in financial applications 6, which hinders reproducibility and makes it difficult for the community to compare results and build on prior works. There is a strong need for an open and collaborative platform where researchers can systematically evaluate their methods under realistic financial conditions. To fill this gap, we initiated the FinRL Contests as a community-driven effort to promote reproducibility, transparency, and benchmarking in financial reinforcement learning.
@@
From 2023 to 2025, we have organized three FinRL contests to support reproducible experimentation and systematic benchmarking. These contests cover diverse tasks such as stock trading, order execution, crypto trading, and LLM-generated signals for FinRL. The contests attracted 200 participants from over 100 institutions and 22 countries in total. They formed 46, 21, and 85 teams across 2023 to 2025, respectively. To ensure transparency and reproducibility, we embrace the open-source community. To support the development of solutions for the FinRL Contest tasks, we provide starter kits for participants, including tutorials, market data, GPU-optimized parallel market environments, and baseline methods. The full list of starter kits and contest websites from 2023 to 2025 is summarized in Table 1. We also encourage participants to open-source their solutions. In addition, we maintain a continuously updated documentation, serving as a detailed guide and resource hub for participants. This documentation includes a detailed FinRL introduction, task descriptions, and baseline solution explanations.
@@
In this paper, we summarize FinRL Contests 2023-2025, which is a series of benchmarking efforts for financial reinforcement learning (FinRL). We detail the formulation of financial tasks as Markov Decision Processes (MDPs), the design of an automated data curation pipeline, the implementation of GPU-optimized parallel market environments, and the establishment of standardized evaluation protocols. In addition, we provide an overview of participant performance and reflect on the organizational experience gained from running these contests.
@@
2025 Github Repo 2025 Website 2025
2024 Github Repo 2024 Website 2024
2023 Github Repo 2023 Website 2023
Documentation FinRL Contests documentation
Table 1: Links to the starter kits and documentation provided for FinRL Contests 2023-2025.
@@ 2.1 Applications of Financial Reinforcement Learning
Financial markets are inherently complex, dynamic, and high-dimensional, making deep reinforcement learning (DRL) a compelling approach for sequential decision-making 5. DRL has been successfully applied to a variety of financial tasks 4, including algorithmic trading, portfolio optimization, and option pricing, where traditional rule-based or supervised methods often fall short in capturing long-term rewards under uncertainty. Beyond conventional reward-driven learning, recent advances in preference-based reinforcement learning offer alternative supervision signals. For example, Direct Preference Optimization (DPO) 11 formulates learning objectives based on pairwise comparisons between trajectories, enabling policy improvement in settings where explicit reward design is difficult or noisy. Such methods hold promise for financial applications that involve subjective goals, human feedback, or implicit utility functions, and may further enhance the robustness and flexibility of DRL in real-world markets.
@@ 2.2 Methods of Financial Reinforcement Learning
A wide range of reinforcement learning methods 3, 4 have been explored for financial decision-making. Among these, algorithms based on policy gradients and actor-critic architectures 12 have received particular attention due to their effectiveness in handling continuous control, high-dimensional action spaces, and non-stationary dynamics. Notable examples include proximal policy optimization (PPO), deep deterministic policy gradient (DDPG), soft actor-critic (SAC), and twin-delayed DDPG (TD3), which are frequently adopted for their robustness and sample efficiency. 13 proposed CVaR Proximal Policy Optimization (CPPO) algorithm, which integrates Conditional Value-at-Risk (CVaR) into its objective based on PPO.
Building on these algorithmic foundations, many studies have explored DRL applications in quantitative finance. For example, Deng et al. 14 propose a recurrent deep reinforcement learning framework that jointly learns market representations and trading policies. Their model demonstrates robust performance across stock and commodity markets by integrating deep feature extraction with sequential decision-making. Avramelou et al. 15 propose a multi-modal embedding approach to integrate both price and sentiment data in DRL-based trading agents. Their method improves trading performance while enabling dynamic adjustment of modality importance without retraining. To address the practical challenges of applying DRL in quantitative trading, Li et al. 16 propose FinRL-Podracer, a scalable cloud-based framework for efficient training and deployment of DRL agents. The system integrates high-performance GPU optimization with automated training workflows, significantly improving both trading performance and development efficiency. While recent studies have advanced DRL applications across various financial domains, systematic evaluation and comparison of different approaches remain an important direction, motivating further efforts toward benchmarking financial reinforcement learning.
@@ 2.3 Benchmarks of Financial Reinforcement Learning
Recent research works have applied deep reinforcement learning (DRL) to quantitative finance 17, 18, 19, 20, 21 by developing customized market environments. Although these open-source libraries provide useful environments and datasets from sources like Yahoo Finance and WRDS, there is still a lack of standardized benchmarks for systematic evaluation. FinRL 1, 2, 22 offers a full pipeline with environments for stock trading, portfolio allocation, and cryptocurrency trading; however, these environments are increasingly insufficient to meet the growing demands of the community. To address the challenges of data quality, survivorship bias, and model overfitting, Liu et al. 9, 10 further introduced FinRL-Meta. FinRL-Meta follows a DataOps paradigm to automatically collect dynamic datasets and construct hundreds of gym-style market environments, reproduces popular DRL trading papers to facilitate research reproducibility, and supports cloud-based deployment and visualization for community-driven performance evaluation. Despite these advances, establishing a unified, large-scale, and competitive benchmark for data-driven financial reinforcement learning remains an open direction, motivating the development of FinRL Contests.
@@ 3 Tasks
Based on FinRL 1, 2, 22, FinRL-Meta 9, 10, and our open-source community, we select several stable and representative trading tasks for the FinRL Contests. In this section, we describe the tasks in the language of Markov Decision Processes (MDPs), discuss the parallelism in the estivate of policy gradients, and describe the specific tasks featured in the FinRL Contests.
@@ 3.1 Tasks in FinRL Contests 2023-2025
We summarize trading tasks of FinRL Contests 2023-2025 in Fig. 1, including daily stock trading with OHLCV data and second-level cryptocurrency trading with LOB data. The stock trading tasks include:
@@
  • FinRL Contest 2023 Data-Centric Stock Trading 1, 2, 22. It is to train a stock trading agent by applying novel data and feature engineering based on market data. It is a multi-asset trading task for 30 constituent stocks in the Dow Jones Index. The task has been clearly defined and well studied in FinRL 1, 2, 22 and FinRL-Meta 9, 10. The task is stable and reproducible, making it ideal for benchmarking in a competition setting. In addition, this task focuses on data and feature engineering, encouraging participants to tackle the challenge of dynamic market data. This aligns with the community’s ongoing interest in exploring dynamic financial data and developing stable multi-asset environments.
@@
  • FinRL Contest 2024 LLM-Engineered Signals with RLMF 23, 24. It is to develop LLMs that can generate actionable trading signals from financial news by using reinforcement learning from market feedback (RLMF). LLMs have been applied in many financial tasks 24, but general-purpose LLMs trained on Internet data cannot capture the intrinsic dynamics of financial markets. To align LLMs with financial markets, this task encourages adapting LLMs using Reinforcement Learning from Market Feedback, as a counterpart of Reinforcement Learning from Human Feedback (RLHF) 25. RLMF utilizes feedback from the financial market as reward signals, enabling LLMs to learn from financial market behaviors. This task was selected for its novelty and potential to push the frontier of LLM-enhanced financial decision-making. It aligns with the growing interest in the FinRL and FinLLM communities in combining FinRL and LLMs.
@@
  • FinRL Contest 2025 FinRL-DeepSeek for Stock Trading 23, 26. It is to develop trading agents by combining FinRL and LLM-generated signals. Unstructured financial texts, such as news and SEC filings, contain valuable information about market sentiment, risk, and events. LLMs, such as DeepSeek-V3, have been applied to extract signals from financial documents 24, which are then integrated into FinRL. This task was selected because it reflects real-world decision-making, where financial analysts and traders rely on both structured market data and unstructured financial text. It also aligns with the growing interest in the FinRL and FinLLM communities in LLM-generated signals and multimodal data exploration.
@@
Different from stock markets, cryptocurrency trading faces greater volatility due to drastic price fluctuations and market sentiment shifts. Cryptocurrency markets operate 24/7, demanding adaptable strategies in a continuous trading environment without traditional market open and close cycles. It is more challenging to develop robust and profitable trading strategies. To reflect these unique challenges, we include cryptocurrency trading tasks in the FinRL Contests. To differentiate from the multi-asset stock trading tasks above, the cryptocurrency trading tasks are designed to trade single-asset Bitcoin at the second level. The cryptocurrency trading tasks include:
@@
  • FinRL Contest 2024 Cryptocurrency Trading with Ensemble Learning 17 is to train a trading agent for Bitcoin by using ensemble methods. RL policies are unstable, whose performance is sensitive to hyperparameters, market noise, and random seeds. Policy instability can also come from value function approximation errors. These issues are amplified in cryptocurrency markets with high volatility. Ensemble methods have shown effectiveness in mitigating policy instability 17. This task extends these ideas to the cryptocurrency trading tasks. In addition, collecting high-quality financial trajectories is often expensive and limited by data access. It causes the sampling bottleneck, especially during training multiple FinRL agents for an ensemble. Therefore, this task encourages participants to employ novel ensemble methods and utilize the provided GPU-optimized parallel environments to improve sampling speed. It also aligns with the need in the financial industry for robust and efficient trading strategies in cryptocurrency markets.
@@
  • FinRL Contest 2025 FinRL-AlphaSeek for Crypto Trading 17 extends the previous task of cryptocurrency trading by introducing an additional focus on factor mining. Factors such as momentum play a critical role in driving trading decisions, enabling traders to design efficient, data-driven strategies. This task encourages participants to develop a two-stage pipeline: (1) factor engineering and selection and (2) ensemble learning to build robust trading agents. It aligns with the industry need for effective alpha and beta signal extraction in cryptocurrency markets.

### v4

@@ Abstract
[Abstract]Financial reinforcement learning (FinRL) is now a practical paradigm for financial engineering. However, applying RL strategies to real-world trading tasks remains a challenge for individuals, as it is error-prone and engineering-heavy. The non-stationarity of financial data, low signal-to-noise ratios, and various market frictions require deep accumulations. Although numerous FinRL methods have been developed for tasks such as stock/crypto trading and portfolio management, the lack of standardized task definitions, real-time high-quality datasets, close-to-real market environments, and robust baselines has hindered consistent reproduction in both open-source community and FinTech industry. To bridge this gap, we organized a series of FinRL Contests from 2023 to 2025, covering a diverse range of financial tasks such as stock trading, order execution, crypto trading, and the use of large language model (LLM)-engineered signals. These contests attracted 200+ participants from 100+ institutions over 20+ countries. To encourage participations, we provided starter kits featuring GPU-optimized parallel market environments, ensemble learning, and comprehensive instructions. In this paper, we summarize these benchmarking efforts, detailing task formulations, data curation pipelines, environment implementations, evaluation protocols, participant performance, and organizational insights. It guides our follow-up FinRL contests, and also provides a reference for FinAI contests alike.
@@
Financial reinforcement learning, FinRL, LLM, benchmark, FinAI, financial engineering
articletype: Article Typejournal: Journalvolume: 00
@@ 1 Introduction
Financial reinforcement learning (FinRL) 1, 2, 3, 4, 5, 6 is an interdisciplinary field that applies reinforcement learning algorithms to financial tasks, such as algorithmic trading, portfolio management, option pricing, hedging, and market making 7, 8, 9, 10, 11. FinRL aims to train trading agents to interact with dynamic financial environments and to optimize their financial decisions autonomously. As FinRL methods continue to advance, two practical challenges have become increasingly apparent. One is policy instability, where small changes in training settings or market environments can lead to large variation in performance. The other is the sampling bottleneck, as collecting high-quality trajectories is often expensive and limited by access to financial data. Additionally, it remains difficult to compare the performance of these RL strategies due to inconsistencies in task definitions, dataset choices, environment implementations, and baselines. These challenges make it difficult to obtain reliable and consistent results, emphasizing the need for reproducible benchmarks that enable fair and systematic evaluation across tasks.
@@
FinRL-Meta 5, 6 partially addressed this issue by transforming financial data into gym-style market environments, which built an automated data curation pipeline to deal with dynamic market data. Such a framework enables different FinRL algorithms to be implemented and evaluated under standardized conditions across various tasks. Still, there is no widely accepted benchmark for evaluating RL agents in financial applications 10, which hinders reproducibility and makes it difficult for the community to compare results and build on prior works. There is a strong need for an open and collaborative platform where researchers can systematically evaluate their methods under realistic financial conditions. To fill this gap, we initiated the FinRL Contests as a community-driven effort to promote reproducibility, transparency, and benchmarking in financial reinforcement learning.
@@
From 2023 to 2025, we have organized a series of FinRL contests to support reproducible experimentation and systematic benchmarking. These contests cover diverse tasks such as stock trading, order execution, crypto trading, and LLM-engineered signals for FinRL. Excluding the ongoing FinAI Contest 2025, the FinRL contests attracted 200+ participants from 100+ institutions over 20+ countries. They formed 46, 21, and 85 teams across 2023 to 2025, respectively. To ensure transparency and reproducibility, all participant teams were encouraged to open-source the solutions. To support the development of solutions for the FinRL Contest tasks, we provide starter kits for participants, including tutorials, market data, GPU-optimized parallel market environments, and baseline methods. The full list of starter kits and contest websites from 2023 to 2025 is summarized in Table 1. In addition, we maintain a continuously updated documentation, serving as a detailed guide and resource hub for participants. This documentation includes a detailed FinRL introduction, task descriptions, and explanations of baseline solutions.
@@
In this paper, we summarize the FinRL Contests 2023-2025 that are our benchmarking efforts for financial reinforcement learning. We detail the formulation of financial tasks as Markov Decision Processes (MDPs), the design of an automated data curation pipeline, the implementation of GPU-optimized parallel market environments, and the establishment of standardized evaluation protocols. In addition, we provide a summary of participant performance and the organizational experience gained from running these contests.
@@
2025 @ IEEE CSCloud* Github Repo 2025 Website 2025
2025 @ IEEE IDS Github Repo 2025 Website 2025
2024 @ ACM 2024 Github Repo 2024 Website 2024
2023 @ ACM 2023 Github Repo 2023 Website 2023
Documentation FinAI Contests documentation
Table 1: Links to the starter kits and documentation provided for FinRL Contests 2023-2025. * The FinAI Contest 2025 @ IEEE CSCloud is ongoing.
@@ 2.1 Applications of Financial Reinforcement Learning
Financial markets are inherently complex, dynamic, and high-dimensional, making deep reinforcement learning (DRL) a compelling approach for sequential decision-making 9. DRL has been successfully applied to a variety of financial tasks 8, including algorithmic trading, portfolio optimization, and option pricing, where traditional rule-based or supervised methods often fall short in capturing long-term rewards under uncertainty. Beyond conventional reward-driven learning, recent advances in preference-based reinforcement learning offer alternative supervision signals. For example, Direct Preference Optimization (DPO) 12 formulates learning objectives based on pairwise comparisons between trajectories, enabling policy improvement in settings where explicit reward design is difficult or noisy. Such methods hold promise for financial applications that involve subjective goals, human feedback, or implicit utility functions, and may further enhance the robustness and flexibility of DRL in real-world markets.
@@ 2.2 Methods of Financial Reinforcement Learning
A wide range of reinforcement learning methods 7, 8 have been explored for financial decision-making. Among these, algorithms based on policy gradients and actor-critic architectures 13 have received particular attention due to their effectiveness in handling continuous control, high-dimensional action spaces, and non-stationary dynamics. Notable examples include proximal policy optimization (PPO), deep deterministic policy gradient (DDPG), soft actor-critic (SAC), and twin-delayed DDPG (TD3), which are frequently adopted for their robustness and sample efficiency. Ying et al. 14 proposed the CVaR Proximal Policy Optimization (CPPO) algorithm, which integrates Conditional Value-at-Risk (CVaR) into its objective based on PPO.
Building on these algorithmic foundations, many studies have explored DRL applications in quantitative finance. For example, Deng et al. 15 propose a recurrent deep reinforcement learning framework that jointly learns market representations and trading policies. Their model demonstrates robust performance across stock and commodity markets by integrating deep feature extraction with sequential decision-making. Avramelou et al. 16 propose a multi-modal embedding approach to integrate both price and sentiment data in DRL-based trading agents. Their method improves trading performance while enabling dynamic adjustment of modality importance without retraining. To address the practical challenges of applying DRL in quantitative trading, Li et al. 4 propose FinRL-Podracer, a scalable cloud-based framework for efficient training and deployment of DRL agents. The system integrates high-performance GPU optimization with automated training workflows, significantly improving both trading performance and development efficiency. While recent studies have advanced DRL applications across various financial domains, systematic evaluation and comparison of different approaches remain an important direction, motivating further efforts toward benchmarking financial reinforcement learning.
@@ 2.3 Benchmarks of Financial Reinforcement Learning
Recent research works have applied deep reinforcement learning (DRL) to quantitative finance 17, 18, 19, 20, 21 by developing customized market environments. Although these open-source libraries provide useful environments and datasets from sources like Yahoo Finance and WRDS, there is still a lack of standardized benchmarks for systematic evaluation. FinRL 1, 2, 3 offers a full pipeline with environments for stock trading, portfolio allocation, and crypto trading; however, these environments are increasingly insufficient to meet the growing demands of the community. To address the challenges of data quality, survivorship bias, and model overfitting, Liu et al. 5, 6 further introduced FinRL-Meta. FinRL-Meta follows a DataOps paradigm to automatically collect dynamic datasets and construct hundreds of gym-style market environments, reproduces popular DRL trading papers to facilitate research reproducibility, and supports cloud-based deployment and visualization for community-driven performance evaluation. Despite these advances, establishing a unified, large-scale, and competitive benchmark for data-driven financial reinforcement learning remains an open direction, motivating the development of FinRL Contests.
@@ 3 Tasks
Based on projects FinRL 1, 2, 3, 4 and FinRL-Meta 5, 6, and feedback from our open-source community, we select several stable and representative trading tasks for the FinRL Contests. In this section, we describe the tasks in the language of Markov Decision Processes (MDPs), discuss the parallelism in the estimate of policy gradients, and describe the specific tasks featured in the FinRL Contests.
@@ 3.1 Tasks in FinRL Contests 2023-2025
We summarize trading tasks of FinRL Contests 2023-2025 in Fig. 1, including daily stock trading with OHLCV data and second-level crypto trading with LOB data. The stock trading tasks include:
@@
  • Data-Centric Stock Trading @ FinRL Contest 2023 1, 2, 3. It is to train a stock trading agent by applying novel data and feature engineering based on market data. It is a multi-asset trading task for 30 constituent stocks in the Dow Jones Index. The task has been clearly defined and well studied in FinRL 1, 2, 3 and FinRL-Meta 5, 6. The task is stable and reproducible, making it ideal for benchmarking in a competition setting. In addition, this task focuses on data and feature engineering, encouraging participants to tackle the challenge of dynamic market data. This aligns with the community’s ongoing interest in exploring dynamic financial data and developing stable multi-asset environments.
@@
  • LLM-Engineered Signals with RLMF @ FinRL Contest 2024 22, 23. It is to develop LLMs that can generate actionable trading signals from financial news by using reinforcement learning from market feedback (RLMF). LLMs have been applied in many financial tasks 23, but general-purpose LLMs trained on Internet data cannot capture the intrinsic dynamics of financial markets. To align LLMs with financial markets, this task encourages adapting LLMs using Reinforcement Learning from Market Feedback, as a counterpart to Reinforcement Learning from Human Feedback (RLHF) 24. RLMF utilizes feedback from the financial market as reward signals, enabling LLMs to learn from financial market behaviors. This task was selected for its novelty and potential to push the frontier of LLM-enhanced financial decision-making. It aligns with the growing interest in the FinRL and FinLLM communities in combining FinRL and LLMs.
@@
  • FinRL-DeepSeek for Stock Trading @ FinRL Contest 2025 22, 25. It is to develop trading agents by combining FinRL and LLM-engineered signals. Unstructured financial texts, such as news and SEC filings, contain valuable information about market sentiment, risk, and events. LLMs, such as DeepSeek-V3, have been applied to extract signals from financial documents 23, which are then integrated into FinRL. This task was selected because it reflects real-world decision-making, where financial analysts and traders rely on both structured market data and unstructured financial text. It also aligns with the growing interest in the FinRL and FinLLM communities in LLM-engineered signals and multimodal data exploration.
@@
Different from stock markets, crypto trading faces greater volatility due to drastic price fluctuations and market sentiment shifts. crypto markets operate 24/7, demanding adaptable strategies in a continuous trading environment without traditional market open and close cycles. It is more challenging to develop robust and profitable trading strategies. To reflect these unique challenges, we include crypto trading tasks in the FinRL Contests. To differentiate from the multi-asset stock trading tasks above, the crypto trading tasks are designed to trade single-asset Bitcoin at the second level. The crypto trading tasks include:
@@
  • crypto Trading with Ensemble Learning @ FinRL Contest 2024 17 is to train a trading agent for Bitcoin by using ensemble methods. RL policies are unstable, whose performance is sensitive to hyperparameters, market noise, and random seeds. Policy instability can also come from value function approximation errors. These issues are amplified in crypto markets with high volatility. Ensemble methods have shown effectiveness in mitigating policy instability 17. This task extends these ideas to the crypto trading tasks. In addition, collecting high-quality financial trajectories is often expensive and limited by data access. It causes the sampling bottleneck, especially during training multiple FinRL agents for an ensemble. Therefore, this task encourages participants to employ novel ensemble methods and utilize the provided GPU-optimized parallel environments to improve sampling speed. It also aligns with the need in the financial industry for robust and efficient trading strategies in crypto markets.
@@
  • FinRL-AlphaSeek for Crypto Trading @ FinRL Contest 2025 17 extends the previous task of crypto trading by introducing an additional focus on factor mining. Factors such as momentum play a critical role in driving trading decisions, enabling traders to design efficient, data-driven strategies. This task encourages participants to develop a two-stage pipeline: 1) factor engineering and selection and 2) ensemble learning to build robust trading agents. It aligns with the industry need for effective alpha and beta signal extraction in crypto markets.
  • FinRL-DeepSeek for Crypto Trading @ FinAI Contest 2025 17 extends the previous two tasks of crypto trading by introducing LLM-engineered signals. crypto markets are highly sensitive to news headlines, tweets, regulatory shifts, and viral narratives. The timely interpretation of market sentiment is critical for crypto trading. LLMs can extract actionable signals from financial news, tweets, and filings. In this task, we encourage participants to explore LLM-engineered signals and integrate them into a FinRL trading agent for crypto trading.